OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2).

OpenStudy (anonymous):

so the equation of the line tells us that the slope of 2x-3y=3 is: y= (-3+2x)/3 y= 2/3 x -1 Therefore slope is: 2/3 the slope of the perpendicular line is the Negative Inverse of the line slope, so that would be: -3/2= m (slope of the perpendicular line) Therefore the equation of the line that is perpendicular to y= 2/3 x -1, is therefore: y= -3/2 x -1

OpenStudy (anonymous):

That' s not in my answer choices.

OpenStudy (anonymous):

what is in ur answer choices?

OpenStudy (anonymous):

y=2/3x-10 y=2/3x+10 y=-3/2x-10 y=-3/2x+10

OpenStudy (anonymous):

oh okey, so this is what you have to do; you have to figure out the value of b. So we got the solve correct, which is m= -3/2 now we need to find b, y=mx+b ==> y=-3/2x+b, Plug in the point (-8,2) to find b. 2=-3/2 * -8 +b -10=b therefore the correct answer is the third choice y=-3/2x-10

OpenStudy (anonymous):

Can you help me with some others please?